OpenStudy (anonymous):

1. What is the product in simplest form? State any restrictions on the variable 4a^5/7b^4 times 2b^2/2a^4 A. 4a^9 / 7b^4, a = 0 , b = 0 B. 4a / 7b^4, a=0 , b=0 C. 7b^2 / 4b^2, a=0 , b=0 D. 4/7a^9 b^6 a=0 b=0

OpenStudy (anonymous):

Product of what,please be more specific.

OpenStudy (anonymous):

4a^5/7b^4 times 2b^2/2a^4

OpenStudy (anonymous):

ok,give me 5 mins and ill have the answer

OpenStudy (anonymous):

okay

OpenStudy (the_fizicx99):

?

OpenStudy (the_fizicx99):

Is that what you meant?

OpenStudy (anonymous):

8a^9/14b^6 i believe.

OpenStudy (anonymous):

no its 4a^5 over 7b^4 times 2b^2 over 2a^4

OpenStudy (anonymous):

and how did you get that?

OpenStudy (anonymous):

times the numbers with the same letters and add the ^ numbers,sorry forget what there called at the moment.

OpenStudy (anonymous):

well thats ot apart of the answer choices hmm

OpenStudy (anonymous):

ok,give me the choices

OpenStudy (anonymous):

okay A. 4a^9 over 7a^6 a=0 , b=0 B. 4a over 7b^2 , a=0 b=0 C. 7b^2 over 4a, a=0 b=0 D. 4/7 a^9 b^6, a=0 b=0

OpenStudy (the_fizicx99):

Oh, quadratic?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Then lets go with A on this one.

OpenStudy (the_fizicx99):

^

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

but give this other guy a chance before you do,he might knw more then i do.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I took the answers i got and cut em in half.

OpenStudy (anonymous):

oh okay haha

OpenStudy (the_fizicx99):

@SolomonZelman <-- he'll know how to guide you :p I'm not this far ahead, sowwie

OpenStudy (solomonzelman):

what is the question again, guys?

OpenStudy (solomonzelman):

that initial question?

OpenStudy (the_fizicx99):

@SolomonZelman its the initial question you nub >.>

OpenStudy (anonymous):

hey,its a valid question.

OpenStudy (anonymous):

4a^5/7b^4 times 2b^2/2a^4

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{4a^5}{7b^4} \times \frac{2b^2}{2a^4} } \]

OpenStudy (solomonzelman):

\[\color{blue}{ \frac{4a^5}{7b^4} \times \frac{2b^2}{2a^4} ~~~->~~~ \frac{4a^5~\times~2b^2}{7b^4~\times~2a^4} ~~~->~~~ \frac{8a^5b^2}{14b^4a^4}. } \] So far so good?

OpenStudy (the_fizicx99):

this is all i see lolll

OpenStudy (anonymous):

lol same here

OpenStudy (solomonzelman):

D a m n it again!

OpenStudy (the_fizicx99):

K fixed mine :3

OpenStudy (solomonzelman):

4a^5 2b^2 8a^5b^2 ---- times ----- ---> -------- 7b^4 2a^4 14a^4b^4 So far so good?

OpenStudy (anonymous):

ok,so i see the step that i missed,much more sense,but the person who asked the question,see what she thinks.

OpenStudy (anonymous):

well its obvious the wrong answer hmm cause that answer isnt in the question at all there is no 8

OpenStudy (solomonzelman):

I am not saying thta's the answer, I am showing you the process and asking if you understand everything I showed so far.

OpenStudy (anonymous):

oh haha i some what understand it

OpenStudy (solomonzelman):

So not entirely?

OpenStudy (anonymous):

no not really :P

OpenStudy (solomonzelman):

I can't teach you now, I am not a good tutor, I can give tips, but teaching staff when I am doing my own work is overwhelming, sorry.

OpenStudy (anonymous):

its alright but do you by chance know the answer :P

OpenStudy (solomonzelman):

I am talking pre-calculus and calculus currently, rockstargirl, but I don;t give out answers like this, this is openSTUDY.

OpenStudy (anonymous):

dam man,ur no fun.